Determine the pH of a 0.045M CaSO3. The Ka for SO3 is 1.3x10^-2
Answer #1

ANSWER : pH = 7.27

Consider reaction, SO 32- (aq) + H2O (l) \rightleftharpoons HSO 3- (aq) + OH - (aq)

For above reaction, K b = [HSO 3- ] [ OH - ] / [SO 32- ] = K w / K a = 10 -14 / 1.3 \times 10 -02 = 7.69 \times 10 -13

Let's Use ICE table.

Concentration (M) SO 32- HSO 3- OH -
Initial 0.045
Change - X +X +X
Equilibrium 0.045 - X X X

Therefore, K b = [HSO 3- ] [ OH - ] / [SO 32- ] = (X) (X) / 0.045 - X = 7.69 \times 10 -13

X 2 / 0.045 - X = 7.69 \times 10 -13

Assume X is very small as compared to 0.045. Therefore, we can write 0.045 - X \simeq 0.045.

X 2 / 0.045 = 7.69 \times 10 -13

X 2 = 0.045 \times 7.69 \times 10 -13

X 2= 3.4605 \times 10 -14

X = 1.8602 \times 10 -07 M = [HSO 3- ] = [ OH - ]

We have relation, [H+ ] [ OH - ] = 10 -14

\therefore [H+ ] = 10 -14 / [ OH - ] = 10 -14 / 1.8602 \times 10 -07 = 5.376 \times 10 -08 M

We have, pH = -log [H+ ] = -log 5.376 \times 10 -08 = 7.27